# Direct Sampling from Conditional Distributions by Sequential Maximum Likelihood Estimations

### v1

@@
2020 Mathematics Subject Classification Numbers. 33C90, 33F99, 62H17, 62R01, 65C05
@@ 1 Introduction
In this paper, we consider the extended log-affine model ([11], Section 4.2.3), whose support may be at the boundary of the simplex. This extension ensures the existence of the maximum likelihood estimate (MLE), whether or not observed counts satisfy constraints to be positive.
@@
$\mathbf{P}(U=u,AU=b)=\frac{|u|!}{\{Z(t,x)\}^{|u|}}\prod_{i\in[d]}t_{i}^{b_{i}}\prod_{j\in[m]}\frac{x_{j}^{u_{j}}}{u_{j}!},$ (2)
where the total number of counts is $|u|:=u_{1}+\cdots+u_{m}$. The conditional distribution given by complete minimal sufficient statistics $b\in\mathbb{N}A:=\sum_{j\in[m]}\mathbb{N}a_{j}$ for $t$ is
$\mathbf{P}(U=u|AU=b)=\frac{1}{Z_{A}(b;x)}\frac{x^{u}}{u!},\quad x^{u}:=\prod_{j\in[m]}x_{j}^{u_{j}},\quad u!:=\prod_{j\in[m]}u_{j}!,$ (3)
where $a_{j}$ denotes the $j$-th column vector of $A$. The support of this distribution is determined by the linear transformation
$\mathcal{F}_{b}(A):=\{v\in\mathbb{N}^{m}:Av=b\},$ (4)
where $\mathbb{N}$ is the set of non-negative integers. The set (4) is called the $b$-fiber associated to the configuration matrix $A$ with the sufficient statistics $b$. The normalization constant, or the partition function
@@
Consider a matrix $A=(a_{ij})\in\mathbb{Z}^{d\times m}$ satisfying the conditions in Definition 1.1 and a vector $b\in\mathbb{N}A$. The Markov lattice $\mathcal{L}_{A}(b)$ is the bounded integer lattice embedded in $\mathbb{N}A$ equipped with the partial order
$\beta\in\mathbb{N}A~{}~{}\text{and}~{}~{}\beta-a_{j}\in\mathbb{N}A\quad\Rightarrow\quad\beta-a_{j}\prec v,$
and the maximum and the minimum are $b$ and 00, respectively.
@@
for the transition from $\beta$ to $\beta-a_{j}$, where $\deg(\beta)$ is the degree of $Z_{A}(\beta;x)$. The cost of the computation of the transition probabilities dominates the computational cost of the direct sampling algorithm. Evaluation of (6) following the definition of the $A$-hypergeometric polynomial (5) is practically impossible since the number of terms in (5) is the cardinality of the fiber (4) that increases rapidly with the total number of counts. In Section 4 of [16], the authors reported some results of numerical experiments of exact computations with rational number arithmetic on a computer algebra system. For two-way contingency tables of the independence model, the direct sampling algorithm demanded a time of about one-tenth time of the Metropolis algorithm when we count the effective sample size of samples taken from the Metropolis chain. On the other hand, for a model without independence, the direct sampling algorithm demanded a thousand times the time the Metropolis algorithm demanded. For the latter case, although the computation is exact, the computational burden would prohibit the practical use of the direct sampling algorithm.
@@
In this paper, we show that the transition probability (6) is the ratio of the uniformly minimum variance unbiased estimator (UMVUE) of the expected count to the total number of counts, where for a vector of observed counts $u\in\mathbb{N}^{m}$, we call $\mu=|u|p$ the vector of expected counts. In the final section of Chapter 1, entitled “Exact Methods,” Haberman [9] argued that the UMVUE was impractical for two main reasons: first, the UMVUE generally does not fall on the model manifold $\mathcal{M}_{A}$ (the toric variety defined by the matrix $A$), and second, it was computationally intensive. In response to this, in the final paragraph of the chapter, she concluded, “Since exact methods are usually impractical, approximate methods must be tried. One such method, that of maximum likelihood, provides estimation and test procedures which involve simple computational methods and which have large sample properties. This method is the subject of the remainder of the monograph.” In this context, the significance of UMVUE seems to have been overlooked, and the discussion of the MLE has dominated the literature.
@@
This paper is structured as follows. In Section 2, we introduce the details of the direct sampling algorithm proposed in [14] and show that the transition probability is the ratio of the UMVUE of the expected count to the total number of counts. Then, we show that the UMVUE and the MLE coincide if and only if a log-linear model is a decomposable graphical model, and the resulting MLE is rational. Lastly, we discuss the use of the MLE and reveal the key how we avoid computation of the connection matrices of the holonomic ideal. Section 3 is devoted to numerical experiments. Using several contingency tables as examples, we demonstrate how the direct sampling algorithm works. Section 4 summarizes the significance of this paper.
@@ 2 Algorithm
where the second last equality holds by $\mathbb{E}[\partial l/\partial\xi_{i}]=0$. The MLEs of the canonical parameters $\hat{\xi}_{i}$, $i\in[d]$ constitute the unique solution of the estimating equations $b_{i}=\partial\psi/\partial\xi_{i}$. Therefore, we have
@@
  • Step 3’: Pick $j_{t}=j$ with probability $\hat{\mu}_{j}(\beta;x)$, $j\in[m]$,
and discuss implementations of Step 2’.
Remark 2.2.
A merit of Algorithm 1 is that we do not have to construct the Markov lattice $\mathcal{L}_{A}(b)$. A Markov transition occurs only between adjacent elements because if otherwise, $Z_{A}(\beta-a_{j};x)=0$ since $\beta-a_{j}\not\in\mathbb{N}A$. For the same reason, a sample path never hits outside of $\mathcal{L}_{A}(b)$, or more precisely, the set $({\rm Cone}(A)\cap\mathbb{Z}^{m})\setminus\mathbb{N}A$, where ${\rm Cone}(A)$ is the polyhedral cone generated by the column vectors of $A$. Note that $Z_{A}(\beta-a_{j};x)=0$ if and only if $\hat{\mu}_{j}(\beta;x)=0$ since the UMVUE and the MLE coincide asymptotically.
@@ 2.1 The UMVUE and Markov transitions
$(\partial^{z^{+}}-\partial^{z^{-}})f=0,\quad v\in{\rm Ker}A\cap{\mathbb{Z}}^{m},\quad\partial^{z}:=\prod_{j\in[m]}\partial_{j}^{z_{j}},$ (9)
@@
These properties gives
@@
$\sum_{j\in[m]}\tilde{\mu}_{j}(b;x)=\sum_{j\in[m]}\frac{x_{j}\partial_{j}Z_{A}(b;x)}{Z_{A}(b;x)}={\deg}(b),$
and
$|u|=\sum_{j\in[m]}\mu_{j}=\mathbb{E}(\sum_{j\in[m]}\tilde{\mu}_{j})={\deg(b)}.$
@@
Before closing this subsection, we mention an observation about the UMVUE. Note that the following proposition does not mean that the UMVUE always coincides with the MLE since the UMVUE generally does not fall on the model manifold $\mathcal{M}_{A}$. If it falls, it should coincide with the MLE by Theorem 2.1. We will be back to this issue in Section 2.2.
@@ 2.2 The UMVUE and the rational MLE
Step 2 of Algorithm 1 is immediate if we know a closed form expression of the MLE in advance. In this respect, we have the following result based on the fact that the number of critical points of the likelihood over the complex space is one if and only if the map from the observed counts to the MLE admits a specific form called the Horn uniformization [10].
@@
Theorem 2.6 means that the time computational complexity of Step 2 of Algorithm 1 is $O(lm)$ if we use this rational MLE. Here, we present a small example in [10, 7] in our context.
@@
where the rows are for $u_{1\cdot}$, $u_{2\cdot}$, $|u|$, $u_{\cdot 1}$, and $u_{\cdot 2}$, respectively. The Markov lattice $\mathcal{L}_{A}(b)$ with $b={}^{t}(1,2,2,1)$ is in Figure 1, and the transition probability (6) becomes
$P(b,b-a_{ij};1)=\frac{\hat{\mu}(b;1)}{|u|}=\frac{u_{i\cdot}}{|u|}\frac{u_{\cdot j}}{|u|},\quad i,j\in[2],$ (14)
@@
Remark 2.9.
The fact that Step 2 of Algorithm 1 is immediate for decomposable graphical models is not surprising, given that we can sample directly by combining urn schemes (see, e.g., [20], Section 9.6). However, an implementation of Algorithm 1 may be easier.
@@
The UMVUE of a log-linear model falls on the model manifold $\mathcal{M}_{A}$ if and only if the model is a decomposable graphical model.
2.3 Approximate algorithm without using connection matrices
@@
For a log-affine model, we can always evaluate the MLE numerically solving the estimating equation (7) by using the iterative proportional scaling (IPS), specifically called the generalized IPS by Darroch and Ratcliff [5]. We rewrite the procedure in our notation. Consider a sample of counts $u\in\mathbb{N}^{m}$ following the log-affine model $\mathcal{M}_{A}$ satisfying $Au=\beta$, where $\beta\in\mathbb{N}A$. For simplicity, we assume $A\in\mathbb{N}^{d\times m}$, $\beta_{i}>0,i\in[d]$, and all columns of $A$ have the same sum denoted by $s$, but a general form is available in [5]. That is
@@
The time computational complexity of Algorithm 1 is $O(lm|u|)$ if we use the rational MLE for Step 2, where $l$ is in Theorem 2.6. For the approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’, the complexity is $O(dm|u|\tau)$ if we use the generalized IPS for Step 2’. Here, (16) defines $\tau$.
@@
$x_{j}Z_{A}(b-a_{j};x)=\tilde{\mu}_{j}(\beta;x)Z_{A}(b;x),\quad j\in[m].$
Comparing this with (18), we see that the first row of the connection matrix $F_{j}(\beta;x)$ can be $(\tilde{\mu}_{j}(\beta;x),0,\ldots,0)$. All we need is an evaluation of the UMVUE $\tilde{\mu}$ exactly or approximately, since the recurrence relation for the first element of the vector $q(\beta;x)$ decouples from the others. Here, the MLE asymptotically coincides with the UMVUE and can be used as an approximation of the UMVUE, but we can compute the MLE without using the recurrence relation.
@@ 3.1 Two-way contingency tables
Then, we assessed the convergence of the Metropolis chain to the stationary distribution. For the Metropolis chain of four pairs of the burn-in and length: $(0,10^{3})$, $(10^{3},10^{4})$, $(10^{4},10^{4})$, and $(10^{4},10^{5})$, respectively, we computed the empirical distribution $\hat{p}_{\rm M}$ of the chi-square values and calculated the total variation distance (20). Here, we set the expected value of each cell in the chi-square to $s$ for both the independence and non-independence models with the odds ratios (21). We show the averages of the results obtained by repeating this procedure 10 times in Table 1. The pair $(0,10^{3})$ gave the poorest results because only a few contingency tables appeared in a Metropolis chain, and they cannot approximate the stationary distribution. The pairs $(10^{3},10^{4})$ and $(10^{4},10^{4})$ were equally good, which implies that the Metropolis chain converged after 10,000 steps. For the pair $(10^{4},10^{5})$, the approximation is even better because the empirical distribution approximates the stationary distribution better by many contingency tables because of the Glivenko–Cantelli theorem. These trends were similar for non-independent models. Another important observation is that for approximate sampling of the non-independence models, the total variation distance did not decrease as $s$ increases. While one would expect the total variation distance to decrease as the bias of the MLE decreases asymptotically, the reverse trend suggests that the Monte Carlo error is relatively large compared to the bias and that the impact of the bias is small in typical settings. Indeed, the total variation distances showed a similar trend for the direct sampling from the independence model.

### v2

@@
2020 Mathematics Subject Classification Numbers. 62R01, 33C90, 33F99, 62H17, 65C05
@@ 1 Introduction
In this paper, we consider the closure, called the extended log-affine model by [11], Section 4.2.3, whose support may be at the boundary of the simplex. This extension ensures the existence of the maximum likelihood estimate (MLE), whether or not observed counts satisfy constraints to be positive.
@@
$\mathbf{P}(U=u,AU=b)=|u|!\prod_{j\in[m]}\frac{p_{j}^{u_{j}}}{u_{j}!}=\frac{|u|!}{\{Z(t,x)\}^{|u|}}\prod_{i\in[d]}t_{i}^{b_{i}}\prod_{j\in[m]}\frac{x_{j}^{u_{j}}}{u_{j}!},$ (2)
where the total number of counts is $|u|:=u_{1}+\cdots+u_{m}$. The conditional distribution given by complete minimal sufficient statistics $b$ in the affine semigroup $\mathbb{N}A:=\{Av:v\in\mathbb{N}^{m}\}$, where $\mathbb{N}$ is the set of non-negative integers, is
$\mathbf{P}(U=u|AU=b)=\frac{1}{Z_{A}(b;x)}\frac{x^{u}}{u!},\quad x^{u}:=\prod_{j\in[m]}x_{j}^{u_{j}},\quad u!:=\prod_{j\in[m]}u_{j}!.$ (3)
The support of this distribution is determined by the linear transformation
$\mathcal{F}_{b}(A):=\{v\in\mathbb{N}^{m}:Av=b\}.$ (4)
The set (4) is called the $b$-fiber associated to the configuration matrix $A$ with the sufficient statistics $b$. The normalization constant, or the partition function
@@
Consider a matrix $A=(a_{ij})\in\mathbb{Z}^{d\times m}$ satisfying the conditions in Definition 1.1 and a vector $b\in\mathbb{N}A$. The Markov lattice $\mathcal{L}_{A}(b)$ is the bounded integer lattice embedded in the affine semigroup $\mathbb{N}A$ equipped with the partial order
$\beta\in\mathbb{N}A~{}~{}\text{and}~{}~{}\beta-a_{j}\in\mathbb{N}A\quad\Rightarrow\quad\beta-a_{j}\prec\beta,$
and the maximum and the minimum are $b$ and 00, respectively. Here, $a_{j}$ denotes the $j$-th column vector of $A$.
@@
for the transition from $\beta$ to $\beta-a_{j}$, where $\deg(\beta)$ is the degree of $Z_{A}(\beta;x)$. The computation of the transition probabilities dominates the computational cost of the direct sampling algorithm. Evaluation of (6) following the definition of the $A$-hypergeometric polynomial (5) is practically impossible since the number of terms in the polynomial is the cardinality of the fiber (4) that increases rapidly with the total number of counts. In Section 4 of [16], the authors reported some results of numerical experiments of exact computations with rational number arithmetic on a computer algebra system. For two-way contingency tables of the independence model, the direct sampling algorithm demanded a time of about one-tenth time of the Metropolis algorithm when we count the effective sample size of samples taken from the Metropolis chain. On the other hand, for a model without independence, the direct sampling algorithm demanded a thousand times the time the Metropolis algorithm demanded. For the latter case, although the computation is exact, the computational burden would prohibit the practical use of the direct sampling algorithm.
@@
In this paper, we show that the transition probability (6) is the ratio of the uniformly minimum variance unbiased estimator (UMVUE) of the expected count to the total number of counts, where for a vector of observed counts $u\in\mathbb{N}^{m}$, we call $\mu:=\mathbf{E}U$ the vector of expected counts. In the final section of Chapter 1, entitled “Exact Methods,” Haberman [9] argued that the UMVUE was impractical for two main reasons: first, the UMVUE generally does not lie in the model manifold $\mathcal{M}_{A}$ (the toric variety defined by the matrix $A$); second, it is computationally intensive. In response to this, in the final paragraph of the chapter, she concluded, “Since exact methods are usually impractical, approximate methods must be tried. One such method, that of maximum likelihood, provides estimation and test procedures which involve simple computational methods and which have large sample properties. This method is the subject of the remainder of the monograph.” In this context, the significance of the UMVUE seems to have been overlooked, and the discussion of the MLE has dominated the literature.
@@
This paper is structured as follows. In Section 2, we introduce the details of the direct sampling algorithm proposed in [14] and show that the transition probability is the ratio of the UMVUE of the expected count to the total number of counts. Then, after summarizing the relevant properties of the MLE, we show that the UMVUE and the MLE coincide if and only if a log-linear model is a decomposable graphical model and the resulting MLE is rational. Lastly, we discuss the use of the MLE and reveal the key how we avoid computation of the connection matrices of the holonomic ideal. Section 3 is devoted to numerical experiments. Using several contingency tables as examples, we demonstrate how the direct sampling algorithm works. Section 4 summarizes the significance of this paper.
@@ 2 Algorithm
Remark 2.1.
A merit of Algorithm 1 is that we do not have to construct the Markov lattice $\mathcal{L}_{A}(b)$. A Markov transition occurs only between adjacent elements because if otherwise, $Z_{A}(\beta-a_{j};x)=0$ since $\beta-a_{j}\not\in\mathbb{N}A$. For the same reason, a sample path never hits outside of $\mathcal{L}_{A}(b)$, or more precisely, the set $({\rm Cone}(A)\cap\mathbb{Z}^{m})\setminus\mathbb{N}A$, where ${\rm Cone}(A):=\{Av:v\in\mathbb{R}^{m}_{\geq 0}\}$ is the polyhedral cone generated by the column vectors of $A$.
2.1 The MLE
@@
where the last equality holds by $\mathbb{E}[\partial l/\partial\xi_{i}]=0$. The MLEs of the canonical parameters $\hat{\xi}_{i}$, $i\in[d]$ constitute the unique solution of the estimating equations $b_{i}=\partial\psi/\partial\xi_{i}$. Therefore, we have
@@
Some components of the MLE become zero if sufficient statistics lie in the boundary of ${\rm Cone}(A)$. Related to Definition 1.2, we have the following proposition.
Proposition 2.3.
Consider a matrix $A=(a_{ij})\in\mathbb{Z}^{d\times m}$ satisfying the conditions in Definition 1.1 and a vector of sufficient statistics $b\in\mathbb{N}A$. If $b-a_{j}\not\in\mathbb{N}A$ for some $j$, then the component of the MLE $\hat{\mu}_{j}(b;x)$ vanishes.
Proof.
Since $b\in\mathbb{N}A$ and $b-a_{j}\not\in\mathbb{N}A$ for some $j$, if we write $b=\sum_{k\in[m]}v_{k}a_{k}$ for some $v\in\mathbb{N}^{m}$ (such $v$ is generally not unique), then $v_{j}=0$. This fact means that $b$ lies in a face of ${\rm Cone}(A)$, or
$b\in\{Av:v\in\mathbb{N}^{m},v_{j}=0\}\subset\{Av:v\in\mathbb{R}^{m}_{\geq 0},v_{j}=0\}.$
On the other hand, $\hat{\mu}$ satisfies (7), which implies $\hat{\mu}_{j}(b;x)=0$. ∎
@@
  • Step 3’: Pick $j_{t}=j$ with probability $\hat{\mu}_{j}(\beta;x)/\nu$, $j\in[m]$,
and discuss the implementation.
@@ 2.2 The UMVUE and Markov transitions
$(\partial^{z^{+}}-\partial^{z^{-}})f=0,\quad z\in{\rm Ker}A\cap{\mathbb{Z}}^{m},\quad\partial^{z}:=\prod_{j\in[m]}\partial_{j}^{z_{j}},$ (9)
@@
These properties give
@@
$|u|=\sum_{j\in[m]}\tilde{\mu}_{j}(b;x)=\sum_{j\in[m]}\frac{x_{j}\partial_{j}Z_{A}(b;x)}{Z_{A}(b;x)}={\deg}(b).$
@@
Before closing this subsection, we mention an observation about the UMVUE. Note that the following proposition does not mean that the UMVUE always coincides with the MLE since the UMVUE generally does not lie in the model manifold $\mathcal{M}_{A}$. If it lies, it should coincide with the MLE by Theorem 2.2. We will be back to this issue in Section 2.3.
@@ 2.3 The UMVUE and the rational MLE
Step 2’ of Algorithm 1 is immediate if we know a closed form expression of the MLE in advance. In this respect, we have the following result based on the fact that the number of critical points of the likelihood over the complex space is one if and only if the map from the observed counts to the MLE admits a specific form called the Horn uniformization [10].
@@
Here, we present a small example in [10, 7] in our context.
@@
where the rows are for $u_{1\cdot}$, $u_{2\cdot}$, $|u|$, $u_{\cdot 1}$, and $u_{\cdot 2}$, respectively. The Markov lattice $\mathcal{L}_{A}(b)$ with $b={}^{t}(1,2,2,1)$ is in Figure 1, and the transition probability of Step 3’ of Algorithm 1 becomes
$\frac{\hat{\mu}(b;1)}{|u|}=\frac{u_{i\cdot}}{|u|}\frac{u_{\cdot j}}{|u|},\quad i,j\in[2],$ (14)
@@
The UMVUE of a log-linear model lies in the model manifold $\mathcal{M}_{A}$ if and only if the model is a decomposable graphical model.
2.4 The approximate algorithm without using connection matrices
@@
Before proceeding, we summarize what we know about Algorithm 1 according to Theorem 2.9 as a proposition.
Proposition 2.11.
For a log-linear model, the approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’ coincides with the original Algorithm 1 and thus becomes exact if and only if the model is a decomposable graphical model.
Theorem 2.7 means that the time computational complexity of Step 2 of Algorithm 1 is $O(lm)$ if we use the rational MLE.
Remark 2.12.
The fact that Step 2 of Algorithm 1 is immediate by the rational MLE for decomposable graphical models is not surprising, given that we can sample directly by combining urn schemes (see, e.g., [20], Section 9.6). However, an implementation of Algorithm 1 may be easier.
Combining Remark 2.1 and Proposition 2.3, we arrive at the following consequence: if sufficient statistics lie in the boundary of the polyhedral cone generated by the column vectors of $A$, the MLE coincides with the UMVUE and vanishes exactly without bias. In practical terms, we obtain
Proposition 2.13.
The approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’ is a Markov chain on the Markov lattice defined in Definition 1.2.
For any log-affine model, we can always evaluate the MLE numerically solving the estimating equation (7) by using the iterative proportional scaling (IPS), specifically called the generalized IPS by Darroch and Ratcliff [5]. We rewrite the procedure in our notation. Consider a sample of counts $u\in\mathbb{N}^{m}$ following the log-affine model $\mathcal{M}_{A}$ satisfying $Au=\beta$, where $\beta\in\mathbb{N}A$. For simplicity, we assume $A\in\mathbb{N}^{d\times m}$, $\beta_{i}>0,i\in[d]$, and all columns of $A$ have the same sum denoted by $s$, but a general form is available in [5]. That is
@@
The time computational complexity of Algorithm 1 is $O(lm|u|)$ if we use the rational MLE for Step 2, where $l$ is in Theorem 2.7. For the approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’, the complexity is $O(dm\tau|u|)$ if we use the generalized IPS for Step 2’. Here, (16) defines $\tau$.
@@
$x_{j}Z_{A}(\beta-a_{j};x)=\tilde{\mu}_{j}(\beta;x)Z_{A}(\beta;x),\quad j\in[m].$
Comparing this with (18), we see that by appropriate basis transformation we can make the first row of the connection matrix to be $(\tilde{\mu}_{j}(\beta;x),0,\ldots,0)$. All we need is an evaluation of the UMVUE $\tilde{\mu}$ exactly or approximately, since the recurrence relation for the first element of the transformed vector decouples from the others. Here, the MLE asymptotically coincides with the UMVUE and can be used as an approximation of the UMVUE, but we can compute the MLE without using the recurrence relation.
@@ 3.1 Two-way contingency tables
Then, we assessed the convergence of the Metropolis chain to the stationary distribution. For the Metropolis chain of four pairs of the burn-in and length: $(0,10^{3})$, $(10^{3},10^{4})$, $(10^{4},10^{4})$, and $(10^{4},10^{5})$, respectively, we computed the empirical distribution $\hat{p}_{\rm M}$ of the chi-square values and calculated the total variation distance (20). Here, we set the expected value of each cell in the chi-square to $s$ for both the independence and non-independence models with the odds ratios (21). We show the averages of the results obtained by repeating this procedure 10 times in Table 1. The pair $(0,10^{3})$ gave the poorest results because only a few contingency tables appeared in a Metropolis chain, and they cannot approximate the stationary distribution. The pairs $(10^{3},10^{4})$ and $(10^{4},10^{4})$ were equally good, which implies that the Metropolis chain converged after 10,000 steps. For the pair $(10^{4},10^{5})$, the approximation is even better because the empirical distribution approximates the stationary distribution better by many contingency tables because of the Glivenko–Cantelli theorem. These trends were similar for non-independence models. Another important observation is that for approximate sampling of the non-independence models, the total variation distance did not decrease as $s$ increases. While one would expect the total variation distance to decrease as the bias of the MLE decreases asymptotically, the reverse trend suggests that the Monte Carlo error is relatively large compared to the bias and that the impact of the bias is small in typical settings. Indeed, the total variation distances showed a similar trend for the direct sampling from the independence model.